Hartley Auto Supply delivers parts to area auto service centers and is replacing its fleet of delivery vehicles. What is the total vehicle life cycle cost of this gasoline engine truck given the information provided in the following​ table? VEHICLE PURCHASE COST ​$22 comma 00022,000 VEHICLE OPERATING COST PER MILE ​$0.180.18 USEFUL LIFE OF VEHICLE 88 years MILES PER YEAR 25 comma 00025,000 MILES PER GALLON 3333 AVERAGE FUEL PRICE PER GALLON ​$2.832.83 The total vehicle life cycle cost of the gasoline engine truck is ​$ nothing. ​(Round your response to the nearest​ dollar.)
Answer #1

Given, Purchase cost = $22000

Useful life = 8 years

Miles per year = 25000

Hence, Total miles = 25000*8 = 200000

Operating cost per mile = $0.18

Hence, Total operating cost = 0.18*200000 = $36000

Miles per gallon = 33

Total gallons of fuel used = 200000/33 = 6060.61 gallons

Fuel price = $2.83 per gallon

Hence, Total fuel cost = 2.83*6060.61 = $17151.52

Hence, Total life cycle cost = Purchase cost + Total Fuel cost + Total Operating cost = 22000 + 17151.52 + 36000 = $75151.52=$75152.